OpenStudy (anonymous):

how to simplify: x/(x+1)

OpenStudy (raden):

is is the integration ?

OpenStudy (raden):

* it

OpenStudy (anonymous):

tried long division and i got (1/(x+1))-1

OpenStudy (anonymous):

no its trying to get in the form (A/(x+1)) + B

OpenStudy (anonymous):

btw the answer is 1-A/(x+1)

OpenStudy (raden):

x/(x+1) cant be simplied again. it already be simplest form but it can reforms like this : x/(x+1) = (x+1-1)/(x+1) = (x+1)/(x+1) + - 1/(x+1) = 1 - 1/(x+1)

OpenStudy (anonymous):

thanks but cant you reform it using long division?

OpenStudy (raden):

we can using long division if the numerator degree is highest from the denomitor